OpenStudy (anonymous):

solve each system using elimination 3p + q = 7 2p - 2q = -6

OpenStudy (anonymous):

To solve by elimination, you multiply an entire equation by LCM to cancel out a variable. In other words, if you want to cancel q, multiply the top equation by 2. Or you could cancel p by multiplying -2 on the top, and positive 3 on the bottom... get it? *2 3p + q = 7 6p + 2q = 14 2p - 2q = -6 2p - 2q = -6 Then you add them together: 6p + 2q = 14 2p - 2q = -6 8p = 8 /8 /8 p = 1 Now you can substitute o into an original equation to find q 3(1) + q = 7 3 + q = 7 -3 -3 q = 4 Then check in both equations: 3(1) + (4) = 7 (check!) 2(1) - 2(4) = -6 (check!)